OpenStudy (anonymous):

find the domain and range of (sin sqrt x)^-2

OpenStudy (anonymous):

\[(\sin \sqrt{x})^{-2}\]

OpenStudy (anonymous):

You could put it in your Ti Calculator in the y= area then press 2nd graph to get a table. It will give you the answer.

OpenStudy (anonymous):

domain is \[x\geq 0\] since you cannot take the square root of a negative number

jimthompson5910 (jim_thompson5910):

plus you'll have a lot of holes because you're dividing by sin(sqrt(x)), so solve sin(sqrt(x)) = 0 to find which values of x to remove

OpenStudy (anonymous):

range is \[[1,\infty)\] since the biggest since can be is 1, and so the smallest \[\frac{1}{\sin^2(\sqrt{x})}\] can be is 1

OpenStudy (anonymous):

oh damn i totally forget about that sorry. yes there an infinite amount of places where the function is not defined

jimthompson5910 (jim_thompson5910):

oh sry solve sin^2(sqrt(x)) = 0 to find the values to remove

OpenStudy (anonymous):

any values for which \[\sin(\sqrt{x})=0\]

OpenStudy (anonymous):

like for example \[\{0,\sqrt{\pi},\sqrt{2\pi},...\}\]